OpenStudy (anonymous):

-1/3 + -1/6 = ? a) 1/3 b) -1/2

OpenStudy (e.mccormick):

Get a common denominator and then add them.

OpenStudy (anonymous):

so the common denominator is 18 right? @e.mccormick

OpenStudy (e.mccormick):

Well, that is a common denominator... but because \(3\times 2 = 6\) I would just use 6. \(\large \ddot \smile\)

OpenStudy (anonymous):

@e.mccormick ok now.

OpenStudy (e.mccormick):

If you use 6, yout don't have to do anything to \(-\dfrac{1}{6}\)

OpenStudy (e.mccormick):

So, what do you do to \(-\dfrac{1}{3}\)?

OpenStudy (anonymous):

@e.mccormick ..

OpenStudy (e.mccormick):

Yes? I am still here. What do you need to change \(-\dfrac{1}{3}\) into so it is over 6?

OpenStudy (anonymous):

I dont know.. @e.mccormick

OpenStudy (anonymous):

@e.mccormick I got the answer 1/3

OpenStudy (anonymous):

I mean -1/2

OpenStudy (e.mccormick):

Well, let me show you something: \(\dfrac{5}{7}\times \dfrac{4}{4}=\dfrac{20}{28}\) Yes, \(-\dfrac{1}{2}\) was it. Let me type that actual problem out in steps using 6 and you might see it.

OpenStudy (e.mccormick):

\(-\dfrac{1}{3}+-\dfrac{1}{6}\) \(-\dfrac{1}{3}\times \dfrac{2}{2}+-\dfrac{1}{6}\) \(-\dfrac{2}{6}+-\dfrac{1}{6}\) \(-\dfrac{3}{6}\) \(-\dfrac{1}{2}\)

OpenStudy (e.mccormick):

See how I just used the fact that 2x3=6 to decide what to do? Because I knew I wanted to make the bottom of the fist fraction into 6, I knew that I had to multiply by a form of 1 that had a 2 in it. Well, \(\tfrac{2}{2}=1\), so I can multily by \(\tfrac{2}{2}\) and it does not change the value of the fraction, it just changes what numbers are used. Once I had 6 on the bottom of both, I added, simplified, and was done.